OpenStudy (anonymous):

Solve the system, using substitution. 4x + y = 4 and y = 4 - 4x A. no solution B. infinitely many solutions I put A is this correct?

OpenStudy (anonymous):

oopsies

OpenStudy (anonymous):

wrong room ^^

OpenStudy (anonymous):

It's ok, Ill answer:P

OpenStudy (anonymous):

I agrre

OpenStudy (anonymous):

\[4x+y=4\ and \ y =-4x+4\] substituting y into \[4x+y=4\] gives you \[4x+(-4x+4)=4\] which would end up with zero because 4x and -4x would cancel and then you would be left with 4=4